Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
derossi, bersanetti, mantovani, spinicelli, nardecchia, palma, masserot - 23:19 Monday 20 November 2023 (62509) Print this report
ISC recovery of carm null 1f (not finished)

After the TCS was switched off (apart from the CH) we procedeed with the recovery of the ITF.

The phases were checked and Diego found that for PRCL a pi had to be added wrt the phase in the central interferometer.

We experienced some unlocks during the engagement of the SSFS (oscillation at 50 Hz) which were cured by tuning the cali SSFS from -3300 to -4000.

Once locked in CARM NULL 1f,  we observed a 2Hz oscillation due to SIB1 ty, which disappeared tuning the PR X gain. At 19:38 UTC we closed the CARM2MC loop. (Regarding this first part of the shift maybe Diego will post a more detailed entry).

Then I tried to adjust the phases for the SR (sr tx b1p dcp phi0 from -0.8--> -0.7 and ty from -1,1 --> -2.5) and the SRCL set loops (os phi from 4.55--> 1.4). However once I closed the SR angular and the SRCL SET loop I saw that the SR TX MAR moved by 1urad and the OS went to -10, so I opened again those loops but it was too late and the ITF unlocked at 21:57 UTC.

I leave the ITF in autorelock carm null 1f (but with both SR angular and SRCL SET loop commented, to be better checked tomorrow).

 

Comments to this report:
mantovani, derossi - 9:50 Tuesday 21 November 2023 (62511) Print this report

This morning, we have tuned the OS phase and engaged (and put in the automation) the SRCL set servo.

derossi, mantovani - 14:49 Tuesday 21 November 2023 (62518) Print this report

We closed the SR angular alignment loop @carm null 1f, which was the last loop to be checked and closed at this step.

We waited for the SRCL SET to bring the OS to 0 and we moved SR TX and TY to find the good alignment, which corresponds to a higher DCP, SRCL SET, and SDB2 B1p DC value. In the aligned position we tuned the phase to bring the I quadrature to 0, we then verified its sign, by moving the SR in angle, and closed the loop (see attached plot).

We updated the metatron .ini and .py

Images attached to this comment
Search Help
×

Warning

×